Circular No. 7104

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SPACE-MIRROR EXPERIMENT On Feb. 4, the Russian Znamya-2.5 space-mirror experiment, attached

to the Progress M-40 spacecraft in low-earth orbit, will be used to illuminate selected nighttime regions of the earth with reflected sunlight at an intensity of up to 10 times that of the full moon. The illumination patch is predicted to be 6 km in diameter and will pass over a given site in ∼ 1 to 2 seconds. Observatories may wish to take precautions to avoid damage to equipment and observations. The experiment is scheduled to last from Feb. 4.482 to 5.092 UT, and announced targets include Kazakstan, Ukraine, Belgium, Germany, Canada, and the northern U.S.A. More details of the

experiment program are available at http://src.space.ru/page 30e.htm
and http://www.skypub.com/news/special/znamya.html. A resolution

adopted by the IAU General Assembly in Kyoto in 1997 notes that the launch of large, luminous objects in orbit around the earth is likely to “have deleterious effects on astronomical observations, [and as] the night sky is the heritage of all humanity, which should therefore be preserved untouched, … the appropriate authorities [should] ensure that the night sky receives no less protection than has been given to the world heritage sites on earth.”

RX J0812.4–3114 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association, reports on behalf of the RXTE ASM team at GSFC and the Massachusetts Institute of Technology: “An analysis of the RXTE All Sky Monitor light curve of the transient x-ray source RX J0812.4–3114 (= LS 992, Motch et al. 1999, A.Ap. 323, 853) obtained between 1996 Jan. 5 and 1999 Jan. 28 reveals a modulation at a period of ∼ 81 days. The source appears to have been in a lower flux state until early in 1998, when the 81-day outbursts became visible at levels of ∼ 1 ASM count/s (10 mCrab). If these outbursts continue, then the next predicted maximum is at ∼ 1999 Mar. 25, and observations are encouraged at that time to search for pulsations from this likely Be/neutron star binary.”

COMET P/1998 U3 (JÄGER) Total visual magnitude estimates: Jan. 5.75 UT, 10.8 (W. Hasubick,

Buchloe, Germany, 0.44-m reflector); 11.02, 10.5 (J. Bortle, Stormville, NY, 0.41-m reflector); 18.91, 10.2 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 27.57, 10.9 (N. Biver, Oahu, HI, 0.26-m reflector).

1999 February 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green